OpenStudy (anonymous):

write the equation of the line passing through (-3,6) and perpendicular to the line whose equations is 6x-y-4=0

OpenStudy (tkhunny):

What is the slope of the line to which it is perpendicular? 6x - y - 4 = 0

OpenStudy (anonymous):

6/1 i think

OpenStudy (tkhunny):

"I think"? You have to do better than that. 6x - y - 4 = 0 y = 6x - 4 The slope is 6, as you have indicated. Why did you hesitate? Knowing that this line has a slope of 6, what it the slope if the line we want? The two lines are perpendicular.

OpenStudy (anonymous):

-6?

OpenStudy (tkhunny):

The two slopes must multiply to -1. 6 * (-6) = -36 -- That's not -1. We need 6 * (Something) = -1 Try again.

OpenStudy (anonymous):

-0.16 then, so do all perpendicular slopes multiply to -1?

OpenStudy (anonymous):

oh wait I think i got it you just flipped the slope, right? sorry for the confusion

OpenStudy (tkhunny):

-0.16 is a rather poor answer. The EXACT answer is -1/6 Not all perpendicular slopes multiply to -1. You must pay attention to vertical and horizontal lines. We can now recast the problem: Write the equation of the line through (-3,6) with a slope of -1/6.

OpenStudy (anonymous):

-3=-1/6 (6) - 2?

OpenStudy (freckles):

The line will have form y=mx+b. You must replace the contracts m and b with numbers. You already said you are replacing my with - 1/6. Are you saying you need to replace b with - 2? Also x and y are variables that means they represent many different values so you leave it as x and y.

OpenStudy (freckles):

m with - 1/6*

OpenStudy (anonymous):

meaning that the new line would look like y=-1/6x-2 ?

OpenStudy (freckles):

Looks good